OpenStudy (anonymous):

In the year 2011, Wali Corporation has 100% payout ratio with earning per share (EPS) of Rs. 12 and required rate of return for such type of investment in the market is 16%. Furthermore, the share price of Wali Corporation with 100% payout is Rs. 75. Now, Management is planning to finance a new project through retained earnings without making any change in current capital structure. For this purpose they have decided to retain 40% in current year and expecting 16% return on it. Required: Whether the change in payout ratio resulted by above planning will influence the share price o

OpenStudy (anonymous):

P=12/0.16=75"NO GROWTH SCENARIO" P=12*0.6/(0.16-0.064)=75,=N.A SOLUTION COZ COMPANY IS INVESTING IN GROWTH ASSET.OR INVESTMENT IN GROWTH ASSETS IN FUTURE. P=12*0.6*(1+D51)/(0.16-0.064)=79.80"INCLUDING GROWTH ASSET"